Solve the system of equations using the elimination method. 4x+10y=2, 3x+5y=5
Answer by checkley75(3666)

4x+10y=2
3x+5y=5 multiply this equation by -2 & add.
-6x-10y=-10
4x+10y=2
--------------
-2x=-8
x=-8/-2
x=4 ans.
3*4+5y=5
12+5y=5
5y=5-12
5y=-7
x=-7/5 ans.
Proof:
4*4+10*-7/5=2
16-70/5=2
16-14=2
2=2